$SPAD/src/input f01qef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
wheret:="s"
m:=5
n:=3
ncolq:=5
lda:=5
zeta:Matrix SF:=
 [[1.2247 ,1.1547 ,1.2649 ]]
a:Matrix SF:=
 [[-4.0 ,-2.0 ,-3.0 ,0.0 ,0.0 ],_
  [0.4085 ,-3.0 ,-2.0 ,0.0 ,0.0 ],_
  [0.3266 ,-0.4619 ,-4.0 ,0.0 ,0.0 ],_
  [0.4082 ,-0.5774 ,0.0 ,0.0 ,0.0 ],_
  [0.2449 ,-0.3464 ,-0.6326 ,0.0 ,0.0 ]]
result:=f01qef(wheret,m,n,ncolq,lda,zeta,a,-1)

———-

References^{[1] nothing}

4